Mardi Received an inheritance of $60,000. She invested part at 7% and deposited the remainder in tax free bonds at 8%. Her total annual income from the investments was $4,300. Find the amount invested at 7%.

If anyone could help that would be great!
thanks.

Answer by checkley71(8403)

.07x+.08(60000-x)=4300 or .07x+4800-.08x=4300 or -.01x=4300-4800
or -.01x=-500 or x=-500/-.01 or x=$50,000 is the the amount invested @ 7%.
PROOF: .07*50000+.08(60000-50000)=4300 or 3500+.08*10000=4300 or
3500+800=4300 or 4300=4300.